# Testing stimulated emission photon directions

### v9

@@ Abstract
While naively laser only causes excitation of external target, e.g. Rabi cycle, STED microscopy or backward ASE in optical amplifiers demonstrate it can also cause its deexitation, however, under uncommon condition of being prepared as excited. These two causalities are governed by absorption-stimulated emission pair of equations, and swap places in perspective of CPT symmetry, however, it means photon direction of stimulated emission should be opposite to usually assumed. This article discusses various arguments and proposes simple direct tests to experimentally verify existence of such backward photon trajectories, complementing consequent forward textbook trajectories. Depending on the results, it could lead to many proposed applications like 2WQC more symmetric quantum computers, or if unsuccessful to macroscopic violation of CPT symmetry, so far probably tested only in microscopic settings.
Keywords: lasers, photonics, STED microscopy, backward ASE, CPT symmetry, ring laser, synchrotron radiation, 2WQC, optical fiber, EDFA amplifier, optical isolator
@@ I Introduction
Figure 1: Top: recent widely commented experiment by Steinberg group [1] has observed medium response (measured phase of 85Rb) before and after the applied impulse, proposing ”photon can spend a negative amount of time in an atom cloud” explanation, which does not seem physical, however, theoretical explanation for this article uses forward/backward-in-time propagators [2].
@@
CPT symmetry (of charge conjugation (C) + parity transformation (P) + time reversal (T)) is shown [4] to be necessary for local Lorentz invariant quantum field theories, hence is generally believed to be satisfied by our nature. It has dozens of successful experimental confirmations [5], however, it seems all tests so far were microscopic - come from observations of single particles. Conscious macroscopic test was not found in literature, important especially due to T time symmetry which might lead to new applications - some are proposed and discussed here. For example through system response before impulse, like recently experimentally observed by Steinberg group as in Fig. 1 [1]111 www.scientificamerican.com/article/evidence-of-negative-time-found-….
@@
Figure 2: Top: CPT symmetry says that if there exists ”laser causes excitation” scenario, also ”laser causes deexciation” should exist. Being CPT symmetry analogs, they should have opposite photon direction and delay sign - we would like to verify it experimentally and apply if successful. For example in STED microscopy diode lasers cause both.
@@
Bottom: proposed STED-like test of this photon direction(s): continuous laser excites fluorescent dye being approximately 3-level medium, and impulse laser causes its deexcitation through stimulated emission - its photon direction can be found from $\Delta t=(d\pm l)/c$ delay between laser impulse and observed reduced intensity by fast photodetector focused on dye spontaneous emission. Textbooks assumption would mean ${}^{\prime}+^{\prime}$ sign here, while CPT symmetry requires ${}^{\prime}-^{\prime}$ sign instead, hence this is also macroscopic test of CPT symmetry, complementing dozes of earlier microscopic tests [5].
@@
Figure 3: In commercially available e.g. erbium-doped optical fiber amplifiers (EDFA, or simpler semiconductor SOA, etc.) [6] there is observed strong backward ASE (amplified spontaneous emission) - increased emission from pumped doped fiber toward laser [7], exactly as in the sought CPT analog of ”laser causes excitation” effect, usually removed by forward optical isolators. Top: Analogously to STED-like test, we can verify this photon direction by measuring delay between the impulse and reduced population of excited atoms of external target, e.g. by monitoring its spontaneous emission. For this purpose, the shown standard forward optical isolator should be removed, or better rotated to backward position. Bottom: simplified setting from [7] seems much more practical for tests: of delay (opposite if CPT analog), separation (leaving forward photons in CPT perspective), transparency (blocked by excited/deexcited?) and potential future applications.
@@
As in Fig. 2, applying CPT symmetry to ”laser causes target excitation” scenario, it becomes ”CPT(laser) causes CPT(target) deexcitation”, replacing absorption equation with stimulated emission, having photons traveling from CPT(target) to CPT(laser), also reversing time delay, what might explain observations of negative time delays, also e.g. [8], finally being a consequence of large number of single emissions/absorptions.
@@
We would like to verify existence of such CPT analog of ”laser causes excitation” scenario with backward photon trajectory, for example in STED-like setting as in Fig. 2. Well known backward ASE in common optical amplifiers, as in Fig. 3, seems exactly the sought scenario, also providing practical realization of such test, and maybe also discussed further potential applications thanks to wide commercial availability of optical fiber hardware.
@@
In contrast, as in Fig. 4, standard textbook view on stimulated emission is energy being kind of knocked forward by incoming photon - with the opposite direction for the emitted photon. There are lots of experimental evidence for existence of such process. While it seems in disagreement with CPT symmetry, bottom of Fig. 4 shows it can be viewed as a consequence of the sought: CPT analog of ”laser causes excitation” scenario.
As there are at least these two possibilities for stimulated emission photon directions, the final judge should be experiment. Related tests were earlier proposed using free electron laser [9] (earlier 2009 post), or ring laser [10]. However, they turned out too difficult to organize. This article simplifies test setting to STED-like, and later to fiber optics amplifiers - determining photon direction through delay. Hence it should be organized soon - to update this article, and depending on the outcome: leading to further discussed novel applications, or indicating macroscopic violation of CPT symmetry.
Absorption/stimulated emission equations are switched in CPT perspective, and e.g. diode laser in STED microscope acts with both on external target (bidirectional beam), also Rabi cycle requires both. There is a natural question if we can separate them - there will be shown arguments that e.g. optical isolator or synchrotron radiation allows to split them - into forward/backward unidirectional beam as in Fig. 6.
Multiple updates of this article were mainly motivated by finding additional indirect experimental evidence in the literature (e.g. for Fig. 1, 3, 5), transparency with textbook view as consequent, convenient optical amplifier realization, and adding multiple potential applications. Also, QInter of our 2WQC group (https://www.qaif.org/2wqc): Vysakh Saji has performed the proposed experiment, obtaining looking confirming results, but without followup. Recently we have started local tests (with Tomasz Urbańczyk and Tomasz Kawalec), which lead to observation/difficulty that most of available laboratory lasers have built-in optical isolator, what seems one of reasons backward ASE properties are practically unknown, also confirmed by discussion with multiple photonics and fiber optics experts. Further tests seem to show pulling of photons as required by CPT symmetry and discussed here - planned to be developed and described in a separate article, also to update this article.
[…]
Center: basic argument for ()($\leftarrow$)( ← ) is CPT symmetry of physics: transforming ”laser causes target excitation” scenario into ”CPT(laser) causes CPT(target) deexcitation”, also switching the stimulated emission-absorption equations, governed by the same (CPT symmetric) Einstein’s coefficients $B_{12}=B_{21}$.
Bottom: schematic explanation that CPT view ()($\leftarrow$)( ← ) also allows to conclude textbook ()($\rightarrow$)( → ) picture, hence both can be true. For photons travelling in ”laser causes excitation/deexcitation” CPT analogous scenarios, we add a medium on the way: looking at transparency in the latter, stimulated emission should fully act on the target, but some photons it produces should be absorbed by medium - returning from CPT perspective it becomes textbook amplification by going through excited medium. All should be coupled by standing waves as in Fig. 7, stimulating/amplifying all these photon emissions.
@@ II STED microscopy based tests
While spontaneous emission photons should have a random direction, the question of concern is direction of stimulated emission photons: ()($\rightarrow$)( → ) forward ano/or ()($\leftarrow$)( ← ) backward? Textbooks assume forward, discussion here based on CPT symmetry suggests backward as fundamental, and forward as its consequence (Fig. 4). Therefore, both can be true, the question to verify is existence of backward ()($\leftarrow$)( ← ). All tests require laser without built-in forward optical isolator.
II-A Delay STED-like test
Assuming there are only these two options: ()($\leftarrow$)( ← ) or ()($\rightarrow$)( → ) and we would like to test existence of the former, the simplest way to distinguish them seems delay. There is literature discussing delay in STED microscopy, like [14, 15, 16], however, data they provide seems insufficient to uniquely answer this question.
@@ II-A Delay STED-like test
Hence (fast) photodetector should be connected to oscilloscope, triggered by depletion laser impulse. Its observed delay $\Delta t$ contains positive contribution $d/c$ for light path from dye to detector, and ±$\pm l/c$ contribution from impulse laser-dye path, of sign for one of the two assumed possibilities - which could be determined from this delay, e.g. from oscilloscope snapshots planned to be included in the next version of this article.
@@
II-B Alternative STED-like tests
Another approach is placing photodetector in depletion laser-dye line behind the dye, allowing to determine ()($\leftarrow$)( ← ) or ()($\rightarrow$)( → ) from intensities. Measuring also angle dependence, we could also test (rather exclude) change of angle hypothesis.
@@ II-B Alternative STED-like tests
Figure 6: Top: ring laser uses optical isolator to enforce unidirectional photon circulation, which would be reversed in CPT perspective - switching the action of absorption and stimulated emission equations, suggesting only one of them should act on external targets - let us refer to them as forward/backward beam. Analogously for just adding optical isolator, or synchrotron sources like wiggler, undulator, free electron laser - reversing electron and photon trajectory in CPT perspective.
@@
Bottom: while CPT symmetry of physics is required by local Lorentz-invariant QFT, e.g. switching forward/backward beams, in theory general relativity allows its realization by travelling through nonorientable wormhole [17], bringing tangible thought experiment valuable for intuitions.
@@
Figure 8: Top: Mathematical analogy between electrodynamics and hydrodynamics (superfluid - no viscosity) from [28].
Center: therefore, also Rabi cycle also needs hydrodynamical analog, tank of liquid with standing wave seems analogous to laser resonator. Being semi-transparent/semi-open in one direction (also in Fig. 7), it should cause left-right swimming of a probe particle outside, through exchange of momentum acting in both directions, using positive/negative pressure. In CPT perspective it looks the same, also with energy/momentum exchange in both directions.
Bottom: while the above bidirectional source is resonator extending to external targets, coupling with them through standing wave e.g. for Rabi oscillations, there are also unidirectional sources. In this analogy, synchrotron radiation is kind of water stream allowing only unidirectional momentum transfer. Optical isolator analog could be some unidirectional valve, analog of ring laser: a pump enforcing unidirectional circulation - with reversed circulation in CPT perspective. Radiation pressure is a vector: can be also negative [29, 30], suggesting to build more symmetric and powerful two-way microfluidic, photonic computers as in Fig. 10.
@@ III CPT symmetry in photonics arguments
Let us gather here some arguments for the expected ()($\leftarrow$)( ← ) possibility mainly based on CPT symmetry. However, as in Fig. 4, it seems to also imply textbook ()($\rightarrow$)( → ) case, hence both can be true - hence practically this is a test of existence of ()($\leftarrow$)( ← ) effect.

### v12

@@ Abstract
While naively laser only causes excitation of external target, e.g. Rabi cycle, STED microscopy or ASE/SASE/SSA demonstrate it can also stimulate its deexitation, however, under uncommon condition of being prepared as excited. These two causalities are governed by absorption-stimulated emission pair of equations, and swap places in perspective of T/CPT symmetry, however, it means photon direction of stimulated emission should be opposite to usually assumed, allowing for negative radiation pressure $\vec{p}=\langle\vec{E}\times\vec{H}\rangle/c$. This article discusses various arguments and proposes simple direct tests to experimentally verify existence of such backward photon trajectories, complementing consequent forward textbook trajectories. Depending on the results, it could lead to many proposed applications like medical, astronomical or 2WQC more symmetric quantum computers. Alternatively, if unsuccessful, it would require macroscopic violation of CPT symmetry, so far tested probably only in microscopic settings.
Keywords: lasers, photonics, STED microscopy, stimulated emission, ASE/SASE/SSA, reverse bias, negative radiation pressure, CPT symmetry, ring laser, synchrotron radiation, 2WQC, optical fiber, EDFA amplifier, optical isolator
@@ I Introduction
Figure 1: Top: recent widely commented experiment by Steinberg group ([1, 2]) has observed medium response (measured phase of 85Rb) before and after the applied impulse, proposing ”photon can spend a negative amount of time in an atom cloud” explanation, which does not seem physical, however, theoretical explanation for this article uses forward/backward-in-time propagators [3].
@@
Figure 2: Top: CPT symmetry says that if there exists ”laser causes excitation” scenario, also ”laser causes deexciation” should exist (needs reversing bias: voltage). Being CPT symmetry analogs, they should have opposite photon direction and delay sign - we would like to verify it experimentally and apply if successful. For example in STED microscopy diode lasers cause both.
@@
Bottom: proposed STED-like test of this photon direction(s): continuous laser excites fluorescent dye being approximately 3-level medium, and impulse laser causes its deexcitation through stimulated emission - its photon direction can be found from $\Delta t=(d\pm l)/c$ delay between laser impulse and observed reduced intensity by fast photodetector focused on dye spontaneous emission. Textbooks assumption would mean $^{\prime}+^{\prime}$ sign here, while CPT symmetry requires $^{\prime}-^{\prime}$ sign instead, hence this is also macroscopic test of CPT symmetry, complementing dozes of earlier microscopic tests [5]. Applying reverse bias for depletion laser as in Fig. 5 brings hope for STED avoiding photobleaching.
Figure 3: Top: swirl-like wave behind marine propeller carries energy, momentum, and angular momentum like photon, also allowing to perform time-symmetry-like transformation by just reversing rotation, getting ”pulling” photons able to pull energy from excited resonator. EM is mathematically close to hydrodynamics (e.g. [6]), suggesting to analogously search for EM ”pulling photons” using e.g. synchrotron radiation: coming from just circulating electron - we could also reverse by just changing electron direction, what should analogously reverse $\vec{p}=\langle\vec{E}\times\vec{H}\rangle/c$ radiation pressure. T/CPT symmetry also switches shown absorption and stimulated emission equations, allowing to estimate strength of the effect. Practically it could be realized e.g. by betatron, synchrotron, free electron laser (FEL) with target positioned for reversed electron trajectory. Bottom: for lower frequencies there might be used analogous to marine propeller e.g. spring-like antenna powered with impulse: reversing $\vec{p}=\langle\vec{E}\times\vec{H}\rangle/c$ radiation pressure for reversed time ($t\to-t,H\to-H$), or used impulse electric potential $V\to-V$ - hopefully allowing both to excite resonator of receiver as in wireless charging, but also speedup it deexcitation/relaxation for reversed impulse.
CPT symmetry (of charge conjugation (C) + parity transformation (P) + time reversal (T)) is shown [7] to be necessary for local Lorentz invariant quantum field theories, hence is generally believed to be satisfied by equations governing our nature, working analogously if running time backward. Moreover, electromagnetism we focus on here, is already T-symmetric, explored e.g. in Wheeler-Feynman absorber theory [8] using retarded and advancing waves. In contrast, 2nd law of thermodynamics is asymmetric, however, it is only statistics of solution we live in, like throwing a rock to lake surface, which is symmetric in equations. Arguments for entropy growth like Boltzmann’s H theorem [9] use mean-field-like approximation called Stosszahlansatz. Applying CPT symmetry before such proof, we would get reversed entropy growth, well summarized in [10] as ”the system is dissipative and decohering in both temporal directions”.
CPT symmetry has dozens of successful experimental confirmations [5], however, it seems all tests so far were microscopic - come from observations of single particles. Conscious macroscopic test was not found in literature, important especially due to T time symmetry which might lead to new applications - some are proposed and discussed here. For example by acting only with stimulated emission on target, negative radiation pressure.
@@
Figure 4: In commercially available e.g. erbium-doped optical fiber amplifiers (EDFA, or simpler semiconductor SOA, etc.) [11] there is observed strong backward ASE (amplified spontaneous emission) - increased emission from pumped doped fiber toward laser [12], exactly as in the sought CPT analog of ”laser causes excitation” effect, usually removed by forward optical isolators. Top: Analogously to STED-like test, we can verify this photon direction by measuring delay between the impulse and reduced population of excited atoms of external target, e.g. by monitoring its spontaneous emission. For this purpose, the shown standard forward optical isolator should be removed, or better rotated to backward position. Bottom: simplified setting from [12] seems much more practical for tests: of delay (opposite if CPT analog), separation (leaving forward photons in CPT perspective), transparency (blocked by excited/deexcited?) and potential future applications. As in Fig. 5, it seems crucial to use reverse bias for the laser, in practice (e.g. [12, 1, 2]) probably included in varying/impulse signal.
@@
Applying CPT symmetry to ”laser causes target excitation” scenario, it becomes ”CPT(laser) causes CPT(target) deexcitation”, replacing absorption equation with stimulated emission, having photons traveling from CPT(target) to CPT(laser), also reversing time delay, what might explain observations of negative time delays e.g. by Steinberg group as in Fig. 1 [1, 2], also system response before impulse like in [13], finally being a consequence of large number of single emissions/absorptions.
@@
We would like to verify existence of such CPT analog of ”laser causes excitation” scenario with backward photon trajectory, for example in STED-like setting as in Fig. 2. Well known backward ASE (amplified spontaneous emission) in common optical amplifiers, as in Fig. 4, seems exactly the sought scenario, also providing practical realization of such test, and maybe also discussed further potential applications thanks to wide commercial availability of optical fiber hardware. Its synchrotron radiation analogs are SASE (self-amplified spontaneous emission) [14] used in FEL [15], and synchrotron self-absorption (SSA) considered in astrophysics [16].
Figure 5: Usually system has tendency to get to the ground state, emitting photon, which e.g. excites some atom, according to absorption equation. However, getting reversed tendency should increase amplitude (probability) of shown symmetric reversed Feynman diagram: causing absorption of photon, emitted by some atom - according to symmetric stimulated emission equation. Beside previous based on synchrotron radiation, diagram shows two examples of obtaining such revered tendency: upper uses membrane GFP (green fluorescent protein) which can act as light-driven protein pump [17], in reversed proton gradient would get tendency to absorb photons, what might be used e.g. by biological organisms for thermoregulation. The bottom is analogously applied reversed bias/potential/electron gradient for light/laser diode, what is also applied for light detectors[18], but should actually also stimulate emission. Impulse laser sources often have e.g. sinc-like shape, so should also include reversed bias, what might explain observations in [1, 2]. Backward ASE [12] probably also uses both forward and reverse bias in varying signal, if successful would be perfect for presented applications. Also Fig. 2, 4 tests should rather use reverse bias. Alternatively, change of $N_{2}$ above/below thermal $N_{2}\sim N_{1}\exp(-\beta(E_{2}-E_{1}))$ of Boltzmann distribution should also increase amplitude of Feynman diagram with emission/absorption there, and absortpion/emission in the target.
@@
In contrast, as in Fig. 7, standard textbook view on stimulated emission is energy being kind of knocked forward by incoming photon - with the opposite direction for the emitted photon. There are lots of experimental evidence for existence of such process. While it seems in disagreement with CPT symmetry, bottom of Fig. 7 shows it can be viewed as a consequence of the sought: CPT analog of ”laser causes excitation” scenario. Radiation pressure seems a good intuition - usually positive stimulating emission outward, but there should be also symmetric: stimulating emission inward, like marine propeller pushing/pulling energy from resonator.
Simple tests with continuous diode laser were performed, but without success. This version of article suggests it was because of using constant forward bias, while as in Fig. 5 it seems crucial to use reverse bias to reverse emission tendency to absorption. Successful realization applied varying signal, impulse lasers instead ([1, 2, 13]), which seem to contain reverse bias contributions. It is planned to be tested soon with reverse bias.
In contrast, promising simpler situations are offered by synchrotron radiation and antennas, as shown in Fig. 3. Synchrotron radiation is emitted by e.g. electron travelling on a circle, being also so from CPT perspective. Related tests were earlier proposed using free electron laser [19] (earlier 2009 post). However, organization of such test has turn out also difficult.
Tests on antennas seem in reach, hopefully to be performed in near future. They are inspired by analogy between electromagnetism and hydrodynamics (e.g. [6], or Barnett effect [20]) - marine propeller is asymmetric, creating both positive and negative pressure, switched by reversing rotation/applying T symmetry - it could remotely excite/deexcite resonator. E.g. spring-like antenna should have analogous asymmetry, switching $\vec{p}=\langle\vec{E}\times\vec{H}\rangle/c$ radiations pressure for T-symmetry reversing magnetic field $H$. We should try to remotely excite resonator this way by applying impulse to such transmitter antenna, and try to speedup its deexcitation/relaxation by applying opposite impulse.
@@ II STED microscopy based tests
While spontaneous emission photons should have a random direction, the question of concern is direction of stimulated emission photons: $(\rightarrow)$ forward ano/or $(\leftarrow)$ backward? Textbooks assume forward, discussion here based on CPT symmetry suggests backward as fundamental, and forward as its consequence (Fig. 7). Therefore, both can be true, the question to verify is existence of backward $(\leftarrow)$. All tests require laser without built-in forward optical isolator, and rather powered by reversed bias.
[…]
Center: basic argument for $(\leftarrow)$ is CPT symmetry of physics: transforming ”laser causes target excitation” scenario into (lasar=) ”CPT(laser) causes CPT(target) deexcitation”, requiring to reverse voltage (bias) as in Fig. 5, switching the stimulated emission-absorption equations, governed by the same (CPT symmetric) Einstein’s coefficients $B_{12}=B_{21}$.
Bottom: schematic explanation that CPT view $(\leftarrow)$ also allows to conclude textbook $(\rightarrow)$ picture, hence both can be true. For photons travelling in CPT analogous scenarios (using reversed bias: voltage), we add a medium on the way: looking at transparency in the latter, stimulated emission should fully act on the target, but some photons it produces should be absorbed by medium - returning from CPT perspective it becomes textbook amplification by going through excited medium. Notice asymmetry for target (also in further Fig. 18): it is now easy to find ground state target, but difficult for excited - making the upper scenario more likely. It is natural to think about it as acting with positive/negative radiation pressure: stimulating outward/inward photon emission (if possible), blocked by deexcited/excited targets on the way.
II-A Delay STED-like test (needs reverse bias)
Assuming there are only these two options: $(\leftarrow)$ or $(\rightarrow)$ and we would like to test existence of the former, the simplest way to distinguish them seems delay. There is literature discussing delay in STED microscopy, like [24, 25, 26], however, data they provide seems insufficient to uniquely answer this question.
@@ II-A Delay STED-like test (needs reverse bias)
Hence (fast) photodetector should be connected to oscilloscope, triggered by depletion laser impulse. Its observed delay $\Delta t$ contains positive contribution $d/c$ for light path from dye to detector, and $\pm l/c$ contribution from impulse laser-dye path, of sign for one of the two assumed possibilities - which could be determined from this delay e.g. with oscilloscope snapshots.
@@
II-B Alternative STED-like tests (needs reverse bias)
Another approach is placing photodetector in depletion laser-dye line behind the dye, allowing to determine $(\leftarrow)$ or $(\rightarrow)$ from intensities. Measuring also angle dependence, we could also test (rather exclude) change of angle hypothesis.
@@ II-B Alternative STED-like tests (needs reverse bias)
Figure 8: Top: ring laser uses optical isolator to enforce unidirectional photon circulation, which would be reversed in CPT perspective for reversed bias - switching the action of absorption and stimulated emission equations, suggesting only one of them should act on external targets - let us refer to them as forward/backward beam. Analogously for just adding optical isolator, or synchrotron sources like synchrotron, betatron, wiggler, undulator, free electron laser, bremsstrahlung - reversing electron and photon trajectory in CPT perspective. FEL reaches narrower spectrum, better coherence, narrower angle [15] - what already seems a consequence of discussed CPT analog of emission: that further electrons stimulate emission from earlier (to be verified).
@@
Bottom: while CPT symmetry of physics is required by local Lorentz-invariant QFT, e.g. switching forward/backward beams, in theory general relativity allows its realization by travelling through nonorientable wormhole [27], bringing tangible thought experiment valuable for intuitions, e.g. switching sign of radiation pressure.
@@
Figure 10: Top: Mathematical analogy between electrodynamics and hydrodynamics (superfluid - no viscosity) from [6]. Center: therefore, also Rabi cycle needs hydrodynamical analog, tank of liquid with standing wave seems analogous to laser resonator. Being semi-transparent/semi-open in one direction (also in Fig. 9), it should cause left-right swimming of a probe particle outside, through exchange of momentum acting in both directions, using positive/negative pressure. In CPT perspective it looks the same, also with energy/momentum exchange in both directions. Another analogy is fan/marine propeller exciting resonator (by swirl-like wave carrying energy+momentum+angular momentum like photon), or deexicitng it if reversing rotation direction. Bottom: while the above bidirectional source is resonator extending to external targets, coupling with them through standing wave e.g. for Rabi oscillations, there are also unidirectional sources. In this analogy, synchrotron radiation is kind of water stream allowing only unidirectional momentum transfer. Optical isolator analog could be some unidirectional valve, analog of ring laser: a pump enforcing unidirectional circulation - with reversed circulation in CPT perspective. Radiation pressure is a vector: can be also negative [38, 39], suggesting to build more symmetric and powerful two-way microfluidic, photonic computers as in Fig. 12.
@@ III CPT symmetry in photonics arguments
Let us gather here some arguments for the expected $(\leftarrow)$ possibility mainly based on CPT symmetry. However, as in Fig. 7, it seems to also imply textbook $(\rightarrow)$ case, hence both can be true - hence practically this is a test of existence of $(\leftarrow)$ effect.